OpenStudy (anonymous):

can someone help me step by step? find the derivative y=ln((e^(7x))/(sqrt(4x-5)))

OpenStudy (anonymous):

before you begin taking the derivative, use the properties of the log to make this expression easier to differentiate

OpenStudy (anonymous):

it is the log of the whole thing right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so first step would be, before beginning to take the derivative, rewrite as \[\ln(e^{7x})-\frac{1}{2}\ln(4x-5)\]

OpenStudy (anonymous):

are those steps clear?

OpenStudy (anonymous):

why is 1/2 there and not (4x-5)^(1/2)

OpenStudy (anonymous):

i used two facts \[\log(\frac{a}{b})=\log(a)-\log(b)\] and \[\log(a^n)=n\log(a)\]

OpenStudy (anonymous):

because \[\log(\sqrt{4x-5})=\log((4x-5)^{\frac{1}{2}})=\frac{1}{2}\log(4x-5)\]

OpenStudy (anonymous):

but 1/2 is outside the () and not inside ?

OpenStudy (anonymous):

You can use the rule like that. As long as it's insside the log.

OpenStudy (anonymous):

the one half comes right out front as a multiplier

OpenStudy (anonymous):

on other words, \(\log(\sqrt{x})=\frac{1}{2}\log(x)\)

OpenStudy (anonymous):

then one more step before differentiating since log and exp are inverse functions, you have \[\log(e^{7x})=7x\]

OpenStudy (anonymous):

okay, so do i need to use product rule for 1/2ln(4x-5)?

OpenStudy (anonymous):

oh no not at all

OpenStudy (anonymous):

\(\frac{1}{2}\) is just a constant , leave it there

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Well you could... But it's a waste of time.

OpenStudy (anonymous):

for example if you wanted the derivative of \(\frac{1}{2}x^3\) you do not use the product rule, you just say \(\frac{3}{2}x^2\)

OpenStudy (anonymous):

gotcha. the power rule^

OpenStudy (anonymous):

so now you have \[7x+\frac{1}{2}\ln(4x-5)\] so the only rule you need now is the chain rule for the second part, and also knowing what the derivative of the log is

OpenStudy (anonymous):

typo there, i meant \[7x-\frac{1}{2}\ln(4x-5)\]

OpenStudy (anonymous):

okay imma try it/

OpenStudy (anonymous):

ok let me know what you get

OpenStudy (anonymous):